# Finite basis Kolmogorov-Arnold networks: domain decomposition for data-driven and physics-informed problems

### v1

@@ 2.4 FBKANs
$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{ic}^{i};\theta^{j})-u(x_{ic}^{i})\right)^{2},$
$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{bc}^{i};\theta^{j})-g(x_{bc}^{i},t_{bc}^{i})\right)^{2},$
$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}\left[\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{r}^{i};\theta^{j})\right]+\mathcal{N}_{x}\left[\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{r}^{i};\theta^{j})\right]\right)^{2},$
$\mathcal{L}_{data}(\theta)$ $=\frac{1}{N_{data}}\sum_{i=1}^{N_{data}}\left(\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x_{data}^{i};\theta^{j})-f(x_{data}^{i})\right)^{2}.$
Here, $\theta=\{\theta^{j}\}_{j=1}^{L}$ is the set of trainable parameters, and the loss function $\mathcal{L}$ contains both a term for data-driven training $\mathcal{L}_{data}$ and terms for physics-informed training $\mathcal{L}_{ic}$, $\mathcal{L}_{bc}$, and $\mathcal{L}_{r}$. In this way, FBKANs are adaptable to given problem characteristics.
For FBKANs, the grid of each local KAN $\mathcal{K}^{j}$ is set separately on the corresponding subdomain $\Omega^{j}$. To do so, we densely sample points $x\in\Omega$ and compute the partition of unity function $\omega_{j}(x)$. We then take the boundaries of each subdomain based on the partition of unity functions as
$a^{j}$ $=\min_{x\text{ s.t. }\omega^{j}(x)>0.0001}x,$
$b^{j}$ $=\max_{x\text{ s.t. }\omega^{j}(x)>0.0001}x.$
This gives a custom grid fit to each subdomain.
We evaluate the models using the relative $\ell_{2}$ error, calculated by
$\frac{||f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x;\theta^{j})||_{2}}{||f(x)||_{2}}.$
Our implementation is based on Jax [65], using the Jax-KAN package [66] for the KAN implementation. For simplicity, we keep the hyperparameters of each KAN on each subdomain fixed, including the width and depth, reducing the total number of hyperparameters. However, in general, the parameters for each individual KAN $\mathcal{K}^{j}$ could be chosen differently.
@@ 3 Data-driven results
3.1 Test 1
In this section, we consider a toy problem
@@ 3.1.1 Scaling of FBKANs
[…]
Figure 4: Results with $N=2,8,$ and $32$ subdomains for eq. 13. (a) Training data and plot of exact $f(x)$. (b) Loss curves (eq. 12). (c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$. (d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$. (e) Scaling results for Test 1 with $L$ subdomains.
@@ 3.1.2 FBKANs with noisy data
One strength of KANs over MLPs is their increased accuracy for noisy data [4]. We now test FBKANs on noisy training data using four subdomains ($L=4$). Therefore, we sample 600 training points from a uniform distribution in $[0,8]$ and evaluate $f(x)$. Then, we add Gaussian white noise with zero mean and varying magnitude to $f(x)$, up to an average relative magnitude of $20\%$. This becomes the training set. The method is then tested on 1 000 evenly spaced points in $[0,8]$ without noise.
The training set and results are shown in fig. 5. In the noisiest case, with relative noise of 18.1% added to the training set, the KAN yields a relative $\ell_{2}$ error of 0.1404, whereas the FBKAN yields a relative error of 0.0646. For all noise levels tested, the FBKAN consistently has a lower relative error than the plain KAN in fig. 5, and the predictions are robust to noisy data.
Figure 5: Results for Test 1 with noisy training data; cf. eq. 13. (a) Example training data and plot of exact $f(x)$ with $9.6\%$ mean relative noise. (b) Loss curves (eq. 12) for an example training with $9.6\%$ mean relative noise. (c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$ with $9.6\%$ mean relative noise. (d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}^{j}(x)$ with $9.6\%$ mean relative noise. (e) Relative $\ell_{2}$ error of the KANs and FBKANs with respect to the magnitude of the noise added to the training data.
3.2 Test 2
We now consider the following equation
@@ 3.2 Test 2
for $(x,y)\in[0,1]\times[0,1].$ This data-driven example exhibits fine-scale oscillations. We test two cases: (1) a KAN and FBKAN with a fixed grid, denoted by KAN-1/FBKAN-1 and (2) a KAN and FBKAN with grid extension, denoted by KAN-2/FBKAN-2. The FBKAN has $L=4$ subdomains. The training set is composed of 10 000 points randomly sampled from a uniform distribution.
In both cases, we begin training with $g=5$ grid points. In the grid extension case, the grid increases every 600 iterations as listed in table 4, and the learning rate drops by $20\%$ each time the grid is updated. As can be observed in fig. 6, the final training loss for FBKAN-1 with a fixed grid is approximately the same as the training loss for KAN-2 with the grid extension approach, even though each FBKAN network has six times fewer grid points. This is also reflected in the relative errors reported in table 1. Comparing the KAN-1 and FBKAN-1 models in fig. 7, KAN-1 struggles to capture the fine-scale features of the data accuratelyand has a larger relative error. As can be seen in fig. 8, KAN-2 and FBKAN-2 both outperform their counterparts with a static grid, but FBKAN-2 is better able to capture the data than KAN-2.
Figure 6: Training loss curves for Test 2.
@@
Table 1: Relative $\ell_{2}$ errors for physics-informed Test 2.
Figure 7: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-1 and FBKAN-1 for Test 2. Note that the error plots have different color scales.
Figure 8: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-2 and FBKAN-2 for Test 2. Note that the error plots have different color scales.
@@ 4.1 Test 3
on $x\in[-4,4]$ with exact solution
@@
We sample 400 collocation points from a uniform distribution ($N_{r}=400$) at every training iteration. The grid starts with $g=5$ and increases by $5$ every 1 000 iterations. The KAN reaches a relative $\ell_{2}$ error of 0.2407 with $g_{final}=20$, whereas the FBKAN for four subdomains ($L=4$) reaches a relative $\ell_{2}$ error of 0.0898 for $g_{final}=20$. For eight subdomains ($L=8$), the FBKAN reaches a relative error of 0.0369 with $g_{final}=15$ and 0.0662 with $g_{final}=10$.
Figure 9: Results for Test 3 with $L=8$. The FBKAN clearly outperforms the KAN, reaching a lower loss with fewer grid points.
4.2 Test 4
We now test the Helmholtz equation
@@ 4.2 Test 4
For each choice of $(a_{1},a_{2})$ we consider three training cases. In the first case, we use higher-order splines with $k=5$ and a fixed grid with $g=5$ (denoted by KAN-1/FBKAN-1). In the second case, we take $k=3$ and use grid extension to increase $g$ by $5$ every $600$ iterations, starting with $g=5$ (denoted by KAN-2/FBKAN-2). We use one hidden layer with width 10 for both cases. Additionally, we test a third case, denoted by KAN-3/FBKAN-3, where we consider the same hyperparameters as KAN-1/FBKAN-1 but with width 5 in the hidden layer. The relative $\ell_{2}$ errors are reported in table 2. In all cases, the FBKAN outperforms the corresponding KAN. The grid extension KAN cases, where we increase the grid size over the training, comes at the expense of increasing the computational time by a factor of 1.4-1.5 in our tests. For $a_{1}=1,a_{2}=4$, the smaller FBKAN-3 network is sufficient to represent the solution accurately. However, for larger values of $a_{1}$ and $a_{2}$, the larger network in FBKAN-1 is necessary. The results for KAN-1/FBKAN-1 are shown in figs. 10 and 11 for $a_{1}=1,a_{2}=4$ and $a_{1}=4,a_{2}=4$, respectively. In fig. 12, we consider $a_{1}=a_{2}=6$ with $L=4,9$ and $16$, demonstrating the further refinement possible with additional subdomains.
@@
Table 2: Relative $\ell_{2}$ errors for physics-informed Test 3.
Figure 10: Results for Test 4 with KAN-1 and FBKAN-1, and $a_{1}=1,a_{2}=4$, $L=4$. The solution along the line $x=0$ is given in the bottom left subfigure. Note that the error plots have different color scales.
Figure 11: Results for Test 4 with KAN-1 and FBKAN-1, and $a_{1}=4,a_{2}=4$, $L=4$. The solution along the line $x=0$ is given in the bottom left subfigure. Note that the error plots have different color scales.
Figure 12: Results for Test 4 with KAN-1 and FBKAN-1 and $a_{1}=a_{2}=6$, with $L=4$ and $L=16$.The solution along the line $x=0$ is given in the middle left subfigure. Note that the error plots have different color scales.
4.3 Test 5
@@
We first consider $c=\sqrt{2}$. The KAN has a relative $\ell_{2}$ error of 0.1402, and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0153, as illustrated in fig. 13. We then consider the harder case with $c=2$, shown in fig. 14. The KAN has a relative $\ell_{2}$ error of 0.1778 and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0587.
Figure 13: Results for Test 5 with $c=\sqrt{2}$ and $L=4$. Note that the error plots have different color scales.
Figure 14: Results for Test 5 with $c=2$ and $L=4.$ Note that the error plots have different color scales.

### v2

@@ 2.4 Finite basis Kolmogorov–Arnold networks
$\mathcal{L}_{ic}(\theta)$ $=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{ic}^{i};\theta^{j})-u(\mathbf{x}_{ic}^{i})\right)^{2},$
$\mathcal{L}_{bc}(\theta)$ $=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{bc}^{i};\theta^{j})-g(\mathbf{x}_{bc}^{i},t_{bc}^{i})\right)^{2},$
$\mathcal{L}_{r}(\theta)$ $=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\frac{\partial}{\partial t}\left[\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{r}^{i};\theta^{j})\right]+\mathcal{N}_{\mathbf{x}}\left[\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{r}^{i};\theta^{j})\right]\right)^{2},$
$\mathcal{L}_{data}(\theta)$ $=\frac{1}{N_{data}}\sum_{i=1}^{N_{data}}\left(\sum_{j=1}^{L}\omega_{j}(\mathbf{x})\mathcal{K}_{j}(\mathbf{x}_{data}^{i};\theta^{j})-f(\mathbf{x}_{data}^{i})\right)^{2},$
using the FBKAN model eq.˜5. Here, $\theta=\{\theta^{j}\}_{j=1}^{L}$ is the set of trainable parameters, and the loss function $\mathcal{L}$ contains both a term for data-driven training $\mathcal{L}_{data}$ and terms for physics-informed training $\mathcal{L}_{ic}$, $\mathcal{L}_{bc}$, and $\mathcal{L}_{r}$. In this way, FBKANs are adaptable to given problem characteristics.
For FBKANs, the grid of each local KAN $\mathcal{K}_{j}$ is defined separately on each subdomain $\Omega^{j}$. To do so, we densely sample points $x\in\Omega$ and compute the partition of unity function $\omega_{j}(x)$. We then take the boundaries of each subdomain based on the partition of unity functions as
$a^{j}$ $=\min_{\begin{subarray}{c}x\text{ such that}\\ \omega^{j}(x)>10^{-4}\end{subarray}}x,$
$b^{j}$ $=\max_{\begin{subarray}{c}x\text{ such that}\\ \omega^{j}(x)>10^{-4}\end{subarray}}x.$
Hence, we obtain a custom grid for each subdomain.
We evaluate the model performance using the relative $\ell_{2}$ error, calculated by
[…]
where $f$ is the ground truth solution.
Our implementation is based on JAX [jax2018github], using the jaxKAN package [Rigas_jaxKAN_A_JAX-based_2024] for the KAN implementation. For simplicity, we employ the same hyperparameters for local KAN models defined on the subdomains, including the width and depth; this reduces the total number of hyperparameters. However, in general, the parameters for each individual KAN $\mathcal{K}_{j}$ could be chosen differently.
@@ 3 Data-driven results
3.1 Data-driven test 1
In this section, we consider a toy problem of fitting the function
@@ 3.1.1 Scaling of FBKANs
First, we consider clean data, $N_{data}=1\,200$, which is sampled uniformly from the interval $[0,8].$ We present results for training on $L=1,2,4,8,16,$ and $32$ subdomains. Note that the function described in eq.˜7 is highly oscillatory, which makes it difficult to capture its behavior with a small KAN. As shown in fig.˜4, increasing the number of subdomains from $L=2$ to $L=8$ and $L=32$ significantly decreases the pointwise error in the FBKAN predictions. Doubling the number of subdomains doubles the number of trainable parameters of the FBKAN model, increasing its expressivity. We observe approximately first order convergence of the relative $\ell_{2}$ error for a large number of subdomains in fig.˜4.
Figure 4: Results with $L=2,8,$ and $32$ subdomains for eq.˜7. (a) Training data and plot of exact $f(x)$. (b) Loss curves (eq.˜6). (c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(x)$. (d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(x)$. (e) Scaling results for Test 1 with $L$ subdomains.
@@ 3.1.2 FBKANs with noisy data
One strength of KANs over MLPs is their increased accuracy for noisy data [liu2024kan]. We now test FBKANs on noisy training data using four subdomains ($L=4$). Therefore, we sample $600$ training points from a uniform distribution in $[0,8]$ and evaluate $f(x)$. Then, we add Gaussian white noise with zero mean and varying magnitude to $f(x)$, up to an average relative magnitude of $20\,\%$, as the training data set. The method is then tested on $1000$ evenly spaced points in $[0,8]$ without noise.
The training set and results are shown in fig.˜5. In the case with the largest noise, with relative noise of 18.1% added to the training set, the KAN yields a relative $\ell_{2}$ error of 0.1404, whereas the FBKAN yields a relative error of 0.0646. For all noise levels tested, the FBKAN consistently has a lower relative error than the plain KAN in fig.˜5, and the predictions are robust to noisy data.
Figure 5: Results for Test 1 with noisy training data; cf. eq.˜7. (a) Example training data and plot of exact $f(x)$ with $9.6\%$ mean relative noise. (b) Loss curves (eq.˜6) for an example training with $9.6\%$ mean relative noise. (c) Plot of the outputs $f(x)$ and $\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(x)$ with $9.6\%$ mean relative noise. (d) Pointwise errors $f(x)-\sum_{j=1}^{L}\omega_{j}(x)\mathcal{K}_{j}(x)$ with $9.6\%$ mean relative noise. (e) Relative $\ell_{2}$ error of the KANs and FBKANs with respect to the magnitude of the noise added to the training data.
3.2 Data-driven test 2
Next, we consider fitting the function
@@ 3.2 Data-driven test 2
for $(x,y)\in[0,1]\times[0,1].$ This data-driven example exhibits fine-scale oscillations. We test two cases: (1) a KAN and FBKAN with a fixed grid, denoted by KAN-1 / FBKAN-1 and (2) a KAN and FBKAN with grid extension, denoted by KAN-2 / FBKAN-2. The FBKAN has $L=4$ subdomains. The training set is composed of 10 00010\,000 points randomly sampled from a uniform distribution.
In both cases, we begin training with $g=5$ grid points. In the grid extension case, the grid increases every $600$ iterations as listed in table˜5, and the learning rate drops by $20\,\%$ each time the grid is updated. As can be observed in fig.˜6, the final training loss for FBKAN-1 with a fixed grid is approximately the same as the training loss for KAN-2 with the grid extension approach, even though each FBKAN network has six times fewer grid points. This is also reflected in the relative errors reported in table˜1. Comparing the KAN-1 and FBKAN-1 models in fig.˜7, KAN-1 struggles to capture the fine-scale features of the data accurately and has a larger relative error. As can be seen in fig.˜8, KAN-2 and FBKAN-2 both outperform their counterparts with a static grid, but FBKAN-2 is better able to capture the data than KAN-2.
Figure 6: Training loss curves for the data-driven test 2.
@@
Table 1: Relative $\ell_{2}$ errors for data-driven test 2.
Figure 7: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-1 and FBKAN-1 for data-driven test 2. Note that the error plots have different color scales.
Figure 8: Reference solution (left), predicted solutions (middle) and relative errors (right) for KAN-2 and FBKAN-2 for data-driven test 2. Note that the error plots have different color scales.
@@ 4.1 Physics-informed test 1
for $x\in[-4,4]$. The analytical solution reads
@@
We sample $400$ collocation points from a uniform distribution ($N_{r}=400$) at every training iteration. We start with a grid with $g=5$ and extend it by $5$ every $1000$ iterations. The KAN reaches a relative $\ell_{2}$ error of 0.2407 with $g_{final}=20$, whereas the FBKAN for four subdomains ($L=4$) reaches a relative $\ell_{2}$ error of 0.0898 for $g_{final}=20$. For eight subdomains ($L=8$), the FBKAN reaches a relative error of 0.0369 with $g_{final}=15$ and 0.0662 with $g_{final}=10$.
Figure 9: Results for physics-informed test 1 with $L=8$. The FBKAN clearly outperforms the KAN, reaching a lower loss with fewer grid points.
4.2 Physics-informed test 2
Next, we test the Helmholtz equation
@@ 4.2 Physics-informed test 2
For each choice of $(a_{1},a_{2})$ we consider three training cases. In the first case, we use higher-order splines with $k=5$ and a fixed grid with $g=5$ (denoted by KAN-1 / FBKAN-1). In the second case, we take $k=3$ and use grid extension to increase $g$ by 5 every $600$ iterations, starting with $g=5$ (denoted by KAN-2 / FBKAN-2). We use one hidden layer with width 10 for both cases. Additionally, we test a third case, denoted by KAN-3 / FBKAN-3, where we consider the same hyperparameters as KAN-1 / FBKAN-1 but with width 5 in the hidden layer. The relative $\ell_{2}$ errors are reported in table˜2.
In all cases, the FBKAN outperforms the corresponding KAN model. The grid extension KAN cases, where we increase the grid size during training, comes at the expense of increasing the computational time by a factor of 1.41.51.4--1.5 in our tests. For $a_{1}=1$ and $a_{2}=4$, the smaller FBKAN-3 network is sufficient to represent the solution accurately. However, for larger values of $a_{1}$ and $a_{2}$, the larger network in FBKAN-1 is necessary. The results for KAN-1 / FBKAN-1 are shown in figs.˜10 and 11 for $(a_{1},a_{2})=(1,4)$ and $(a_{1},a_{2})=(4,4)$, respectively. In fig.˜12, we consider $a_{1}=a_{2}=6$ with $L=4,9$ and $16$, demonstrating the further refinement possible with additional subdomains.
@@
Table 2: Relative $\ell_{2}$ errors for physics-informed test 2.
Figure 10: Results for physics-informed test 2 with KAN-1 and FBKAN-1, and $a_{1}=1,a_{2}=4$, $L=4$. The solution along the line $x=0$ is given in the bottom left subfigure. Note that the error plots have different color scales.
Figure 11: Results for physics-informed test 2 with KAN-1 and FBKAN-1, and $a_{1}=4,a_{2}=4$, $L=4$. The solution along the line $x=0$ is given in the bottom left subfigure. Note that the error plots have different color scales.
Figure 12: Results for physics-informed test 2 with KAN-1 and FBKAN-1 and $a_{1}=a_{2}=6$, with $L=4$ and $L=16$.The solution along the line $x=0$ is given in the middle left subfigure. Note that the error plots have different color scales.
4.3 Physics-informed test 3
@@
For $c=\sqrt{2}$, the KAN model has a relative $\ell_{2}$ error of 0.1402, and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0153, as illustrated in fig.˜13. We then consider the more challenging case with $c=2$, shown in fig.˜14. The KAN has a relative $\ell_{2}$ error of 0.1778 and the FBKAN with $L=4$ has a relative $\ell_{2}$ error of 0.0587.
Figure 13: Results for physics-informed test 3 with $c=\sqrt{2}$ and $L=4$. Note that the error plots have different color scales.
Figure 14: Results for physics-informed test 3 with $c=2$ and $L=4.$ Note that the error plots have different color scales.
5 Multilevel FBKANs
Multilevel FBPINNs (MLFBPINNs) were introduced in [dolean_multilevel_2024] to extend FBPINNs to include multiple levels of domain decomposition simultaneously. The motivation is to allow for approximating components on various different scales. Notably, it had been observed that low-frequency components, also denoted coarse components, cannot be learned well with the one-level approach if the numbers of subdomains are being increased; this makes it necessary to add one or multiple coarser corrections; cf. [dolean_finite_2024, dolean_multilevel_2024]. In MLFBPINNs, the FBPINN architecture in eq.˜4 is modified to accommodate $N$ levels of domain decomposition with each level $l$ having an overlapping domain decomposition with $J^{(l)}$ subdomains
$D^{(l)}=\left\{\Omega_{j}^{(l)}\right\}_{j=1}^{J^{(l)}}.$
We take partition of unity functions $\omega_{j}^{(l)}$. As in the single level case, we then require
[…]
The MLFBPINN architecture is given by
$f_{\theta}(\mathbf{x})=\frac{1}{N}\sum_{l=1}^{N}\sum_{j=1}^{J^{(l)}}\omega_{j}^{(l)}(x)f_{j}^{(l)}(\mathbf{x};\theta_{j}^{(l)}),$ (10)
where again $f_{j}^{(l)}(\cdot;\theta_{j}^{(l)})$ is the neural network with parameters $\theta_{j}^{(l)}$ that corresponds to the subdomain $\Omega_{j}^{(l)}$ in level $l$. The extension to KANs, resulting in MLFBKANs, is then straightforward:
$f(\mathbf{x})\approx\frac{1}{N}\sum_{l=1}^{N}\sum_{j=1}^{J^{(l)}}\omega_{j}^{(l)}(x)\mathcal{K}_{j}^{(l)}(\mathbf{x};\theta^{j}),$
where $\mathcal{K}_{j}^{(l)}$ is the KAN corresponding to the subdomain $\Omega_{j}^{(l)}$ on level $l$.
In this section, we show that MLFBKANs can improve training over FBKANs for solutions with higher or multiple frequencies by considering the Helmholtz equation from section˜4.2 and a two-dimensional multiscale Laplacian problem inspired by [dolean_multilevel_2024]. In order to indicate how many subdomains on the different levels are being used we employ the notation $L=n_{1},n_{2},\ldots$; it means that $n_{1}$ subdomains are used on the first level, $n_{2}$ on the second level, and so on.
5.1 Multilevel physics-informed test 1
We first consider the Helmholtz equation from section˜4.2, choosing larger values of $a_{1}=a_{2}$ to increase the computational intensity of training the MLFBKAN. We consider $a_{1}=a_{2}=8$ and $a_{1}=a_{2}=10$. Error results are reported in table˜3. We compare the results across different architectures in figs.˜15 and 16. In both cases, the MLFBKAN with three or four levels outperforms the single level FBKAN, even with 36 subdomains. Indeed, for $a_{1}=a_{2}=8$, even the MLFBKAN with two levels outperforms the single level FBKANs. These results suggest that MLFBKANs can be more accurate than FBKANs for highly oscillatory problems, while reducing the total number of subdomains needed for accurate results.
Figure 15: Results for multilevel physics-informed test 1 with $a_{1}=1,a_{2}=8$. Note that the error plots have different color scales.
Figure 16: Results for multilevel physics-informed test 1 with $a_{1}=1,a_{2}=10$. Note that the error plots have different color scales.
5.2 Multilevel physics-informed test 2
Finally, we consider a multiscale Laplacian problem also employed in [dolean_multilevel_2024]. We take
$-\nabla^{2}u(x,y)$ $=f(x,y),$ $(x,y)\in[0,1]\times[0,1],$
$u(0,y)$ $=u(1,y)=0,$ $y\in[0,1],$
$u(x,0)$ $=u(x,1)=0,$ $x\in[0,1],$
with
$f(x,y)=\frac{2}{M}\sum_{i=1}^{M}(2^{i}\pi)^{2}\sin(2^{i}\pi x)\sin(2^{i}\pi y).$
The exact solution is then given by
$u(x,y)=\frac{1}{M}\sum_{i=1}^{M}\sin(2^{i}\pi x)\sin(2^{i}\pi y).$
From table˜3, for the case of $M=5$, the MLFBKANs offer a significant reduction of error over the FBKANs. These predictions and errors are shown in fig.˜17. The errors indicate that the solution can be accurately captured with the finest resolution domain decomposition having 16 subdomains.
Figure 17: Results for multilevel physics-informed test 2 with $M=5$. Note that the error plots have different color scales.
$a_{1}=a_{2}=8$ $a_{1}=a_{2}=10$ $M=5$
KAN 0.89089 0.92729 1.08556
FBKAN, L=4 0.91962 0.87100 0.33474
FBKAN, L=16 0.28707 0.72251 0.04175
FBKAN, L=36 0.33562 0.51766 0.09343
MLFBKAN, N = 2, L=1, 4 0.24593 0.87294 0.10461
MLFBKAN, N = 3, L=1, 4, 16 0.06353 0.27380 0.02926
MLFBKAN, N = 4, L=1, 4, 16, 36 0.04231 0.12012 0.03066
Table 3: Relative $\ell_{2}$ errors for the multilevel physics-informed test case 1 (middle columns) and test case 2 (right column).